# EnsemW2S: Can an Ensemble of LLMs be Leveraged to Obtain a Stronger LLM?

### v1

@@ F.1 Different rating for all the datasets
F.2 Comparison between probability based combination with logit based combination of the tokens, during generation and evaluation of combined weak experts.
Figure 11: This figure shows a comparison between probability based combination with logit based combination of the tokens.
@@
F.3 Comparison between different window lengths for "sample and token weighing".
Figure 12: This figure shows a comparison between different token windows for pythia 70m model.
Figure 13: This figure shows a comparison between different token windows for pythia 410m model.
@@
F.4 Supervised-Fine Tuning task for Quartz Question-Answer Dataset
@@
Weak Model Performance Strong Model Performance W2S Performance Weak Model Performance Strong Model Performance W2S Performance
@@
70M 160M 0.4987 0.5115 0.5077 0.5255 3.5% 0.4668 0.4847 0.48 0.4898 2%
70M 410M 0.4987 0.602 0.5077 0.5255 3.5% 0.4885 0.4949 0.4643 0.4923 6%
70M 1B 0.4987 0.6276 0.5051 0.5255 4% 0.486 0.4758 0.4758 0.5026 5.6%
70M 1.4B 0.4987 0.6926 0.5026 0.5153 2.5% 0.486 0.4974 0.4719 0.5089 7.8%
70M 2.8B 0.4987 0.7551 0.5 0.5115 2.3% 0.4872 0.5 0.4923 0.4987 1.3%
160M 410M 0.5013 0.6008 0.5 0.5281 5.6% 0.4694 0.4923 0.4758 0.4834 1.6%
160M 1B 0.5013 0.6403 0.5026 0.523 4.1% 0.4872 0.5 0.4681 0.4898 4.6%
160M 1.4B 0.5013 0.713 0.5077 0.5217 2.8% 0.4668 0.5051 0.4936 0.5038 2.1%
160M 2.8B 0.5013 0.7117 0.5077 0.5153 1.5% 0.4847 0.4847 0.4949 0.5128 3.6%
410M 1B 0.5689 0.6403 0.551 0.551 0% 0.4936 0.4911 0.4921 0.5179 5.2%
410M 1.4B 0.5778 0.6926 0.5816 0.5918 1.8% 0.4936 0.4949 0.5268 0.537 1.9%
410M 2.8B 0.5561 0.7551 0.5599 0.611 9.1% 0.5013 0.5115 0.5434 0.5485 0.9%
1B 1.4B 0.6084 0.6888 0.5982 0.6288 5.1% 0.5472 0.4923 0.5536 0.574 3.7%
1B 2.8B 0.6197 0.7398 0.6288 0.6543 4.1% 0.5497 0.5026 0.5855 0.5957 1.7%
1.4B 2.8B 0.699 0.7538 0.6926 0.713 2.9% 0.588 0.477 0.6161 0.6288 2.1%
@@
F.4.1 Supervised-Fine Tuning task for ARC Question-Answer Dataset
@@
Figure 16: ARC Dataset (Random): This figure shows bar plots comparing accuracy values of weak model performance, w2s model performance (baseline and ours) and strong model performance (oracle) for one specific run of experiments. Values are also mentioned in table 8.
@@
Figure 17: ARC Dataset (Easy-Hard): This figure shows bar plots comparing accuracy values of weak model performance, w2s model performance (baseline and ours) and strong model performance (oracle) for one specific run of experiments. Values are also mentioned in table 8.
Weak Model Size (Pythia) Strong Model Size (Pythia) Data Separation: Random Improv (%) Data Separation: Easy-Hard Improv (%)
Weak Model Performance Strong Model Performance W2S Performance Weak Model Performance Strong Model Performance W2S Performance
Baseline Ours Baseline Ours
70M 160M 0.2381 0.2526 0.2457 0.244 -0.7 0.16 0.221 0.2201 0.2244 2
70M 410M 0.2509 0.2867 0.2688 0.273 1.6 0.16 0.1894 0.1672 0.1783 6.6
70M 1B 0.2381 0.3114 0.2858 0.2875 0.6 0.16 0.1937 0.1962 0.2048 4.4
70M 1.4B 0.2381 0.3166 0.2927 0.3003 2.6 0.16 0.2124 0.1945 0.2133 9.7
70M 2.8B 0.2483 0.3524 0.3106 0.3208 3.3 0.16 0.2671 0.2159 0.2321 7.5
160M 410M 0.2423 0.291 0.2816 0.285 1.2 0.175 0.1826 0.1792 0.1834 2.3
160M 1B 0.2423 0.3157 0.2782 0.2858 2.7 0.175 0.215 0.1945 0.2048 5.3
160M 1.4B 0.2312 0.3234 0.3038 0.3166 4.2 0.175 0.2287 0.2082 0.2125 2.1
160M 2.8B 0.2423 0.3541 0.3089 0.3268 5.8 0.175 0.2679 0.2218 0.2338 5.4
410M 1B 0.2739 0.3114 0.2884 0.2935 1.8 0.1993 0.2133 0.2005 0.2031 1.3
410M 1.4B 0.2739 0.3328 0.3148 0.3225 2.4 0.1993 0.2227 0.209 0.2176 4.1
410M 2.8B 0.2739 0.3643 0.3183 0.3225 1.3 0.1993 0.2654 0.227 0.2415 6.4
1B 1.4B 0.3003 0.337 0.3029 0.3174 4.8 0.2108 0.2227 0.2142 0.2167 1.2
1B 2.8B 0.3012 0.3481 0.3217 0.3259 1.3 0.1954 0.2594 0.2355 0.2372 0.7
1.4B 2.8B 0.3119 0.3481 0.3148 0.3294 4.6 0.2125 0.2517 0.2304 0.2457 6.6
Table 8: This table shows weak to strong generalization using random as well as easy-hard data-splits for ARC dataset. As compared to previous tables 6 and 7, here we run experiment once and note the improvement of our method with respect to the baseline.
Appendix G Broader Impact

### v2

@@ H.2 Difficulty rating for all the datasets
H.3 [Ablation Study: ] Comparison between probability based combination with logit based combination of the tokens, during generation and evaluation of combined weak experts.
Figure 9: This figure compares probability-based combination with logit-based combination of the tokens across different AdaBoost rounds. Here we show improvement from the baseline where baseline is single model. The orange bars represent logit-based combination, while the blue bars represent probability-based combination, showing that probability-based combination performs better.
@@
H.4 [Ablation Study: ] Comparison between different window lengths for “sample and token weighing”.
Figure 10: This figure compares different token window lengths for the Pythia 70M model across various AdaBoost rounds. The plots show improvements over the baseline, where the baseline represents a single model. The different bars (red, blue, green, yellow, and pink) correspond to window lengths of 1, 2, 4, 8, and 16, respectively. We observe that, overall, all window lengths perform similarly. Window length in EmsemW2S plays a role only during sampling step.
Figure 11: This figure compares different token window lengths for the Pythia 410M model across various AdaBoost rounds. The plots show improvements/decline over the baseline, where the baseline represents a single model. Thus, the black colored bars show decline. The different bars (red, blue, green, yellow, and pink) correspond to window lengths of 1, 2, 4, 8, and 16, respectively. We observe that, overall, all window lengths perform similarly. Window length in EmsemW2S plays a role only during sampling step.
Appendix I Cost Analysis of EmsemW2S
Training Cost of Weak Learners:
Each weak learner is trained sequentially, as its performance is contingent upon the outputs of the preceding weak learner. Consequently, while the GPU load may be lower because each weak learner is small as compared to the collection of weak-learners, the overall training time is directly proportional to the number of weak learners utilized.
This is because the input and output token count for each weak learner during training remains approximately constant, as suggested by Adaboost. Only the frequency of samples are adjusted based on weights. In EnsemW2S we sample the tokens by token-weights but eventually combine the sampled tokens while masking the ones not sampled, thus keeping the total tokens approximately similar and training time for each weak-learner independent of the tokens sampled. In the practical superalignment case, pre-trained weak learners can be used, which may further mitigate concerns regarding sequentially training the weak learners.
Strong Model Training and Inference:
The strong model is trained using labels generated by the weak learners and is evaluated on standard datasets. Therefore, the training cost and inference cost associated with the strong model remains unchanged.
Inference Cost of Weak Learners:
The generation process can be executed in parallel as well as sequentially, resulting in a GPU load for generation or clock time for generation respectively, that scales linearly with the number of weak learners. For decoding, once the token-level distributions generated by the weak learners are combined using EmsemW2S algorithm, efficient decoding algorithms like speculative decoding can be employed to produce the final response. However, this is not the focus of this work.
We do agree that the primary increase in computational cost compared to the baseline—where a single weak expert supervises the student—comes from multi-LLM decoding. However, its important to consider that unlike other multi-LLM works Anonymous (2025); Du et al. (2023), we do not use an additional reward or judge model for combining multi-LLM outputs, significantly reducing hardware demands. Moreover, we rely on the same labeled data rather than collecting new data to train separate models. For this same reason we use datasets which do not have very long outputs and are option based.
Appendix J Results
J.1 Baselines for weak-expert performance comparison
In this section we show one additional baseline to compare weak model’s performance. Basically instead of using adaptive weights of the model we use fixed model weights i.e. $\alpha_{t}=1\forall t=[1,T]$. This is similar to bagging. You can find the results in the table 2.
Qwen-1.5B Qwen-3B
Random (ID) Easy-Hard (OOD) Random (ID) Easy-Hard (OOD)
Single Model 5 Models Single Model 5 Models Single Model 5 Models Single Model 5 Models
ARC 46.02 47.00 38.14 38.22 54.02 52.75 42.92 43.00
Quartz 74.23 74.36 72.83 71.55 81.25 76.02 80.99 76.02
Math 47.17 47.20 45.42 46.00 61.74 56.3 50.16 50.02
Table 2: In this table we compare performance of weak experts on additional baseline where we combine all the model (specifically 5 models) by giving each of them equal weight. The column 5 models depicts that. We also mention the baseline shown in the main table which uses the first model. We observe that for all the case the 5-model baseline perform worse or similar to the single model baseline.
J.2 Results on ARC Dataset for generation task
J.3 Results on Sciq Dataset for Binary Classification Task
Figure 13: Binary Classification Task. Top figure shows a bar plot comparing w2s generalization of our method (grey) with a baseline (blue) from Burns et al. (2023) using accuracy values(%) for different combinations of weak and strong model pairs for random data split (top bar-plot) and easy-hard split(bottom bar-plot). Bottom figure shows a line plot comparing the accuracy and performance gap recovered values (PGR). The left two figures are for random data split, while the right two figures are for the easy-hard split to show e2h generalization.
Scaling Law. In Figure 13 (line plot), we see less PGR recovery for the Qwen-1.8B model even though it is similar in size to GPT-2-xl. Similarly, in the bar plot, we see a drastic difference between the oracle performance of GPT2xl and Qwen-1.8B. This is because the Qwen models series are more capable even after being the same size. Thus, model size is not a good metric, but model capability is a better metric for differentiating between weak and strong models.
Appendix K Aggregated plots
Figure 14: Aggregated results for Quartz Data on Generation Task and Sciq Data on Binary Classification Task for both random and easy-hard data splits. We aggregate results for three experimental runs will different seed across all model pairs similar to Burns et al. (2023).
Appendix L Generative Task Details
@@ Appendix I Cost Analysis of EmsemW2S
L.1 Supervised-Fine Tuning task for Quartz Question-Answer Dataset
@@
W2S Performance W2S Performance
@@
Pythia-70M Pythia-160M 0.5077 0.5255 3.5% 0.48 0.4898 2%
Pythia-70M Pythia-410M 0.5077 0.5255 3.5% 0.4643 0.4923 6%
Pythia-70M Pythia-1B 0.5051 0.5255 4% 0.4758 0.5026 5.6%
Pythia-70M Pythia-1.4B 0.5026 0.5153 2.5% 0.4719 0.5089 7.8%
Pythia-70M Pythia-2.8B 0.5 0.5115 2.3% 0.4923 0.4987 1.3%
Pythia-160M Pythia-410M 0.5 0.5281 5.6% 0.4758 0.4834 1.6%
Pythia-160M Pythia-1B 0.5026 0.523 4.1% 0.4681 0.4898 4.6%
Pythia-160M Pythia-1.4B 0.5077 0.5217 2.8% 0.4936 0.5038 2.1%
Pythia-160M Pythia-2.8B 0.5077 0.5153 1.5% 0.4949 0.5128 3.6%
Pythia-410M Pythia-1B 0.551 0.551 0% 0.4921 0.5179 5.2%
Pythia-410M Pythia-1.4B 0.5816 0.5918 1.8% 0.5268 0.537 1.9%
Pythia-410M Pythia-2.8B 0.5599 0.611 9.1% 0.5434 0.5485 0.9%
Pythia-1B Pythia-1.4B 0.5982 0.6288 5.1% 0.5536 0.574 3.7%
Pythia-1B Pythia-2.8B 0.6288 0.6543 4.1% 0.5855 0.5957 1.7%
Pythia-1.4B Pythia-2.8B 0.6926 0.713 2.9% 0.6161 0.6288 2.1%
@@
L.1.1 Supervised-Fine Tuning task for ARC Question-Answer Dataset
@@
Figure 17: ARC Dataset (Random): This figure shows bar plots comparing accuracy values of weak model performance, w2s model performance (baseline and ours) and strong model performance (oracle) for one specific run of experiments. Values are also mentioned in table 7 and 6.
@@
Figure 18: ARC Dataset (Easy-Hard): This figure shows bar plots comparing accuracy values of weak model performance, w2s model performance (baseline and ours) and strong model performance (oracle) for one specific run of experiments. Values are also mentioned in table 7 and 6.